OpenStudy (anonymous):

Explain how to graph x – 2y = 4 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

slope = (y2 - y1)/(x2 - x1) where (x1,y1) and (x2,y2) are the two points. It doesn't matter which point is (x1,y1) and which is (x2,y2). slope = (55 - 35)/(150 - 10) = 20/140 = 1/7 Note that likewise, slope = (35 - 55)/(10 - 150) = -20/(-140) = 1/7

OpenStudy (anonymous):

(55-35)/(150-10)=20/140= 1/7. The slope is 1/7

OpenStudy (anonymous):

thanks heres a medal